Circular No. 6179

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

PKS 1622–297 J. R. Mattox, University of Maryland and Universities Space Research

Association; S. Wagner, Landessternwarte Heidelberg; T. A. McGlynn, Compton Observatory Science Support Center; M. Malkan, University of California, Los Angeles; J. F. Schachter, Harvard-Smithsonian Center for Astrophysics; and P. Sreekumar, Goddard Space Flight Center, on behalf of the the Compton Observatory EGRET Team, communicate: “We report the EGRET detection of quasar PKS 1622–297. The flux at E > 100 MeV during June 6–11 was at least an order of magnitude higher than upper limits obtained in previous observations. The EGRET observation is scheduled to continue until July 5. Observations at other wavelengths may be of great value. This object is in the night sky, but only 17^{o} from the galactic center. The NASA/IPAC Extragalactic Database gives an optical magnitude of 20.5 (and extinction of 1.00 B magnitude) and a 31-GHz flux of 2.2 Jy. Contact mattox@astro.umd.edu for more information.”

JUPITER I (IO) E. Lellouch, Observatoire de Paris, Meudon; G. Paubert and R.

Moreno, Institut de Radio Astronomie Millimétrique (IRAM), Granada; D. F. Strobel, The Johns Hopkins University; and M. Belton, National Optical Astronomy Observatories, report: “We have observed Jupiter’s satellite Io with the IRAM 30-m radiotelescope at Pico Veleta (Spain) on May 30 and 31, and June 22–24, detecting the SO (6,5)–(5-4) transition at 219.949 GHz on both the leading and the trailing sides of Io. The line properties are similar to those of the SO_{2} lines observed since 1990 with the same telescope (i.e., an antenna temperature line contrast of ∼ 60 mK and a FWHM width of ∼ 0.65 MHz). We have also obtained a likely detection of the weaker SO (4,3)–(3-2) line at 138.179 GHz, with a line contrast of ∼ 20 mK. These observations represent the first detection of SO in the atmosphere of Io, and a preliminary interpretation suggests an SO/SO_{2} relative abundance of a few percent.”

COMET 6P/D’ARREST Total magnitude estimates: June 19.96 UT, 16.0 (H. Mikuž, Ljubl-

jana, Slovenia, 0.36-m reflector + CCD + V filter); 24.42, 13.1 (A. Hale, Cloudcroft, NM, 0.41-m reflector, visual); 27.03, 13.6 (Mikuž).

1995 June 27 Daniel W. E. Green